Break-Even Analysis

Media outlets often have websites that provide in-depth coverage of news and events. Portions of these websites are restricted to members who pay a monthly subscription to gain access to exclusive news and commentary.

These websites typically offer a free trial period to introduce viewers to the website. Assume that during a recent fiscal year, one outlet spent $6,578,470 on a promotional campaign for its website that offered two free months of service for new subscribers. In addition, assume the following information:

Number of months an average new customer stays with the service
(including the two free months)

25 months
Revenue per month per customer subscription $32
Variable cost per month per customer subscription $11

Determine the number of new customer accounts needed to break even on the cost of the promotional campaign. In forming your answer, (1) treat the cost of the promotional campaign as a fixed cost, and (2) treat the revenue less variable cost per account for the subscription period as the unit contribution margin.
accounts
Answer #1

Break even accounts = fixed cost / contribution margin per account

Per account revenue = $32* 23 months (2months are free)

=736$ per account.

Variable cost per account = $11*25 months.

275$

Contribution margin per unit = revenue per account-variable cost per account

=736-275

461$

Breakeven accounts = 6578470/461=14270

New customer accounts needed to breakeven om cost of promotional campaign=14270 accounts